pleasePassed Failed 20 16 Results from a civil servant exam are shown in the table to the right. Is there sufficient evidence to su find p-value
Answer #1
#observed frequency
passed failed Total
White candidates 20 16 36
Minoritoy Candidates 8 28 36
total 28 44 72

#Expected frequency

Ei=Σrow*Σcolumn/Σtotal passed failed Total
White candidates 14 22 36
Minoritoy Candidates 14 22 36
total 28 44 72
chi square    χ2 =(Oi-Ei)2/Ei passed failed Total
White candidates 2.571428571 1.636363636 4.207792
Minoritoy Candidates 2.571428571 1.636363636 4.207792
total 5.142857143 3.272727273 8.415584

\chi^2=\sum{}(Oi-Ei)2/Ei=8.41558

#test statistics=8.416

degree of freedom =(r-1)*(c-1)=(2-1)*(2-1)=1

P-value=P(\chi^2>8.416)=0.0037

value of chi-square obtain from chi-square table with the corresponding degree of freedom

#P-value=0.0037